OpenStudy (anonymous):

solve the following quadratic equation by factoring 9x-5x^2=-2

OpenStudy (anonymous):

Use complete square

OpenStudy (anonymous):

9x-5x^2+2=0 -5x^2+9x+2=0 5x^2-9x-2=0 5x^2-10x+x-2=0 5x(x-2)-1(x-2)=0 (5x-1)(x-2)=0 x=1/5 ,2